# HSR-Enhanced Sparse Attention Acceleration

### v1

@@
Bo Chen bc7b@mtmail.mtsu.edu. Middle Tennessee State University.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhizhou Sha shazz20@mails.tsinghua.edu.cn. Tsinghua University.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.
Large Language Models (LLMs) have demonstrated remarkable capabilities across various applications, but their performance on long-context tasks is often limited by the computational complexity of attention mechanisms. This paper introduces a novel approach to accelerate attention computation in LLMs, particularly for long-context scenarios. We leverage the inherent sparsity within attention mechanisms, both in conventional Softmax attention and ReLU attention (with $\mathsf{ReLU}^{\alpha}$ activation, $\alpha\in\mathbb{N}_{+}$), to significantly reduce the running time complexity. Our method employs a Half-Space Reporting (HSR) data structure to rapidly identify non-zero or “massively activated” entries in the attention matrix. We present theoretical analyses for two key scenarios: attention generation and full attention computation with long input context. Our approach achieves a running time of $O(mn^{4/5})$ significantly faster than the naive approach $O(mn)$ for attention generation, where $n$ is the context length, $m$ is the query length, and $d$ is the hidden dimension. We can also reduce the running time of full attention computation from $O(mn)$ to $O(mn^{1-1/\lfloor d/2\rfloor}+mn^{4/5})$. Importantly, our method introduces no error for ReLU attention and only provably negligible error for Softmax attention, where the latter is supported by our empirical validation. This work represents a significant step towards enabling efficient long-context processing in LLMs, potentially broadening their applicability across various domains.
@@ 1 Introduction
In practical LLM applications, there are two scenarios for attention computation depending on the context length $n$ and query length $m$. The first case, $m=\Theta(1)$, represents the iterative text generation based on the pre-computed Key Value Cache (KV), which stores the intermediate attention key and value matrices. The second case, $m=\Theta(n)$, represents the full self-attention computation before text generation or the cross-attention computation. However, in both cases, when the context window $n$ becomes larger, the running time will increase correspondingly, i.e., it will be linear and quadratic in $n$ for $m=\Theta(1)$ and $m=\Theta(n)$, respectively. Thus, reducing the running time of attention computations with long context input becomes essential to minimize response latency and increase throughput for LLM API calls.
In this work, we introduce novel methods to reduce the running time complexity for both cases, i.e., $m=\Theta(1)$ and $m=\Theta(n)$. Our approach is inspired by the inherent sparsity found within attention mechanisms. Numerous prior studies have highlighted the significant sparsity in the attention matrix [13, 4, 50, 70, 61]. This manifestation of sparsity in Softmax attention is that a large number of attention scores, i.e., $QK^{\top}$, concentrate on a small number of entries, which is known as “massive activation”. Due to this nature, Softmax attention can be accelerated by only calculating the entries that contain large attention scores, introducing negligible approximation errors [85, 38].
Moreover, when considering ReLU attention (with $\mathsf{ReLU}^{\alpha}$ activation, $\alpha\in\mathbb{N}_{+}$), we can accelerate the attention computation without any approximation error. ReLU attention is another attention mechanism used in transformer architecture, substituting the conventional Softmax activation function with ReLU, which has demonstrated performance comparable to Softmax attention in various downstream tasks [75, 26]; see Section 2 for more details. In the following, we present the formal definition of ReLU attention.
@@
To expedite the computation, the critical task is to identify the large/non-zero entries for Softmax/ReLU attention, respectively. To do so, we utilize the half-space reporting (HSR) data structure, which is introduced in [1] to address the half-space range reporting problem. This is a fundamental problem in computational geometry and can be formally defined as follows:
Definition 1.3 (Half-space range reporting [1, 66]).
In our framework, we define the half-space as the region where the attention scores (the inner products of key and query vectors) exceed some threshold. We leverage this data structure to expedite the identification of non-zero entries within the ReLU attention matrix and large entries in Softmax attention. Consequently, we can compute the ReLU attention only based on those non-zero entries without any approximation error, and compute the Softmax attention based on entries larger than threshold with negligible approximation errors, resulting in a substantial reduction in computation time. When $m=\Theta(1)$, our methods can significantly accelerate ReLU and Softmax attention computation time over the naive approach from $O(mn)$ to $O(mn^{4/5})$ with pre-processed KV cache. When $m=\Theta(n)$, our online methods can also accelerate ReLU and Softmax attention computation time over the naive approach from $O(mn)$ to $O(mn^{1-1/\lfloor d/2\rfloor}+mn^{4/5})$. In more details, when $m=\Theta(1)$ and for any $d\in\mathbb{N}_{+}$, our Algorithm 2 can achieve the fast generation with pre-processed KV cache in $O(mn^{4/5})$ (Theorem 4.1 and Theorem 4.2)). When $m=\Theta(n)$, our Algorithm 3 can achieve the full attention computation in $O(mn^{1-1/\lfloor d/2\rfloor}+mn^{4/5})$ including HSR initialization time and query time (Theorem 5.1 and Theorem 5.2). Thus, our methods can improve both the generation speed and full attention computation for long input context, i.e., $n$ being excessively large. Furthermore, our empirical results in Section 7 show that the approximation error associated with Softmax attention utilizing “massive activated” entries only is small in practice, which is consistent with our theoretical analysis.
@@
  • To the best of our knowledge, this is the first work incorporating the HSR data structure with attention computation, to reduce the running time complexity with the help of the sparsity within the attention mechanisms.
  • Theoretically, we provide rigorous proofs for reducing the computational time (1) for ReLU attention generation from $O(mn)$ to $O(mn^{4/5})$ (Algorithm 2 and Theorem 4.1); (2) for full ReLU attention computation from $O(mn)$ to $O(mn^{1-1/\lfloor d/2\rfloor}+mn^{4/5})$ (Algorithm 3 and Theorem 5.1), without incurring any approximation error in both cases.
  • We achieve the same running time speed up for the conventional Softmax attention, and we give rigorous theoretical proofs to ensure that the resulting approximation error remains negligible (Theorem 4.2, 5.2 and Theorem 4.3).
@@
Roadmap. Section 2 presents related work. In Section 3, we introduce essential concepts and key definitions used this paper. In Section 4, we present our main results, i.e., guarantees on run time reduction and approximation error. In Section 5, we introduce the extension of our method on full attention computation. In Section 6, we provide a brief summary of the techniques used in our proof. In Section 7, we provide our empirical results of evaluating three mainstream LLMs with Softmax attention with top-$r$ indices on different $r$. In Section 8, we discuss the potential of extending our method to other activation functions. In Section 9, we concludes our algorithm and contributions.
@@ 2 Related Work
Attention acceleration for long context input. Long context window is essential for transformer based LLMs in many downstream tasks. However, due to the quadratic time complexity associated with self-attention mechanisms, transformers are usually hard to inference efficiently. Numerous methods have been proposed to enhance the inference efficiency. One approach involves using alternative architectures as proxies for attention to support faster inference, such as Mamba [21, 17], PolySketchFormer [34], Hopfield Models [29, 74, 28, 79, 73, 24, 25] and Linearizing Transformers [81, 53]. However, the broad applicability of these methods across different applications and modalities remains to be fully validated. Another line of research focuses on approximating attention matrix computation [5, 6, 7, 27, 83, 48, 56, 16, 46, 40, 23, 18, 45]. Nevertheless, these methods often rely on assumptions that may not be practical. For instance, some approaches use polynomial methods to approximate the exponential function, which requires all entries to be bounded by a small constant. However, our HSR-enhanced attention framework is designed based on practical observation and validated by empirical support. These advancements in attention mechanisms and transformer efficiency not only improve general model performance but also play a crucial role in enhancing in-context learning capabilities, where models leverage information from the immediate context to perform tasks without fine-tuning. We refer the readers to some other related works [14, 44, 39, 41, 47, 48, 49, 64, 80, 63, 42, 67, 59, 65, 82, 68, 67].
@@
ReLU attention. ReLU attention is an innovative mechanism that employs the ReLU activation function in place of the traditional Softmax function for attention computation. Previous studies have highlighted the promise potential of ReLU attention in various domains. From empirical side, [75] has demonstrated that incorporating ReLU as the activation function in vision transformers enhances performance on downstream tasks. [62] has shown that transformers equipped with ReLU attention outperform those with Softmax attention, particularly when dealing with large key-value memory in machine translation tasks. From theoretical side, the scale-invariant property of ReLU attention [37] facilitates the scalability of transformer networks. Furthermore, [9, 20] have shown that the inherent properties of ReLU attention contribute positively to the learning process of transformer models. Another key advantage of ReLU attention is that the ReLU function effectively sets all negative values to zero, allowing us to bypass these non-contributory elements during attention computation, thereby reducing the running time of attention computation. Importantly, omitting these zero and negative entries does not introduce any error into the final output of the ReLU attention mechanism.
@@
Half-space reporting (HSR) data structure. The Half-Space Reporting (HSR) data structure, initially proposed by [1], was developed to address the half-space range reporting problem. The expedited range query capability inherent to HSR has been demonstrated to significantly enhance computational efficiency across a variety of tasks, as evidenced by numerous previous works in the literature. Studies such as [32] and [11] have applied HSR to facilitate solving general linear programming (LP) problems. Another line of research has highlighted HSR’s potential in expediting the training process of contemporary neural networks [58, 22]. There is also a collection of research that concentrates on leveraging HSR for the advancement of solutions to geometric and graphical challenges [15, 31, 19].
@@ 3 Preliminary
In Section 3.1, we introduce notations used in the paper. In Section 3.2, we introduce a modified version of Softmax attention that operates on a specific subset of indices. It defines the top-$r$ nearest neighbors Softmax attention, which focuses on the most relevant entries in the attention matrix. In Section 3.3, we describe the massive activation property for attention mechanisms. In Section 3.4, we present a data structure for efficiently solving the half-space range reporting problem.
@@ 3.1 Notations
Here, we introduce basic notations used in this paper. For any positive integer $n$, we use $[n]$ to denote set $\{1,2,\cdots,n\}$. We use $\operatorname*{\mathrm{Var}}[]$ to denote the variance. For two vectors $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $X_{i,j}$ to denote the $i$-row, $j$-th column of $X\in\mathbb{R}^{m\times n}$. We use $\|A\|_{\infty}$ to denote the $\ell_{\infty}$ norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{\infty}:=\max_{i\in[n],j\in[d]}|A_{i,j}|$.
3.2 Softmax Attention with Index Set
Definition 3.1 (Input with index set).
In this work, we consider calculating the Softmax attention on the “massive activation” index set, where we define the “massive activation” index set as the top-$r$ indices. We introduce our definition for top-$r$ indices of Softmax attention as follows:
Definition 3.2 ( Top-$r$ indices Softmax attention ).
Let $q\in\mathbb{R}^{d}$, $K,V\in\mathbb{R}^{n\times d}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ denote the indices of top-$r$ entries of $qK$, where $|\mathsf{NN}(r,q,K)|=r$. Let $\widehat{K},\widehat{V}\in\mathbb{R}^{r\times d}$ and $\overline{K},\overline{V}\in\mathbb{R}^{(n-r)\times d}$ be defined in Definition 3.1. We define the top-$r$ nearest neighbors (NN) Softmax attention computation $\widehat{\mathsf{Attn}}_{s}(q,K,V)\in\mathbb{R}^{d}$ as follows:
where
In Definition 3.2, we view the “massive activated” entries as the top-$r$ entries. Therefore, we only calculate the Softmax attention based on $\widehat{K},\widehat{V}\in\mathbb{R}^{r\times d}$, instead of $K,V\in\mathbb{R}^{n\times d}$.
3.3 Massive Activation
Definition 3.3 (Massive activation property).
Let $\gamma\in[0,1]$, $\beta_{1}\geq\beta_{2}\geq 0$. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ denote the indices of top-$r$ entries of $qK$. We define $(\gamma,\beta_{1},\beta_{2})$ massive activation for a query $q\in\mathbb{R}^{d}$ and key cache $K\in\mathbb{R}^{n\times d}$, if the following conditions hold:
An intuitive understanding of Definition 3.3 is that, the summation of “massive activated” entries dominates the summation of all entries, and the entries we ignored only contributes little to the final summation. Therefore, it is reasonable for us to omit those non “massive activated” entries.
Remark 3.4.
There are many distributions satisfying the property in Definition 3.3, such as (1) $K$ drawing from any subexponential distribution, e.g., multivariate Laplace distributions, (2) $K$ drawing from any mixture of Gaussian distribution with $n^{1-\gamma}$ Gaussian clusters.
@@ 3.4 Half-Space Reporting (HSR) Data Structure
Algorithm 1 Half Space Report Data Structure
1:data structure HalfSpaceReport
4:end data structure
We restate the result from [1] for solving the half-space range reporting problem. The interface of their algortihm can be summarized as in Algorithm 1. Intuitively, the data-structure recursively partitions the set $S$ and organizes the points in a tree data-structure. Then for a given query $(a,b)$, all $k$ points of $S$ with $\operatorname{sgn}(\langle a,x\rangle-b)\geq 0$ are reported quickly. Note that the query $(a,b)$ here defines the half-space $H$ in Definition 1.3. We summarize the time complexity of HSR data structure as follows:
Corollary 3.5 (HSR data-structure time complexity [1], informal version of Corollary A.7).
Let ${\cal T}_{\mathsf{init}}$ denote the pre-processing time to build the data structure, ${\cal T}_{\mathsf{query}}$ denote the time per query and ${\cal T}_{\mathsf{update}}$ time per update. Given a set of $n$ points in $\mathbb{R}^{d}$, the half-space range reporting problem can be solved with the following performances:
@@
4 Main Results on Attention Generation
In this section, we present our key findings regarding attention generation, $m=\Theta(1)$, for both ReLU and Softmax attention mechanisms. Across both scenarios, we have reduced the time complexity from a naive $O(mn)$ to $O(mn^{4/5})$. Specifically, for the ReLU attention model, we have managed to accelerate the processing time without introducing any approximation errors. In the case of Softmax attention, our technique results in only an insignificant approximation error.
Algorithm 2 Attention generation
1:data structure AttentionGeneration $\triangleright$ Lemma 6.2

### v2

@@
Bo Chen bc7b@mtmail.mtsu.edu. Middle Tennessee State University.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhizhou Sha shazz20@mails.tsinghua.edu.cn. Tsinghua University.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the UC, Berkeley.
Large Language Models (LLMs) have demonstrated remarkable capabilities across various applications, but their performance on long-context tasks is often limited by the computational complexity of attention mechanisms. We introduce a novel approach to accelerate attention computation in LLMs, particularly for long-context scenarios. We leverage the inherent sparsity within attention mechanisms, both in conventional Softmax attention and ReLU attention (with $\mathsf{ReLU}^{\alpha}$ activation, $\alpha\in\mathbb{N}_{+}$), to significantly reduce the running time complexity. Our method employs a Half-Space Reporting (HSR) data structure to identify non-zero or “massively activated” entries in the attention matrix. We present theoretical analyses for two key scenarios: generation decoding and prompt prefilling. Our approach achieves a running time of $O(mn^{4/5})$ significantly faster than the naive approach $O(mn)$ for generation decoding, where $n$ is the context length, $m$ is the query length, and $d$ is the hidden dimension. We can also reduce the running time for prompt prefilling from $O(mn)$ to $O(mn^{1-1/\lfloor d/2\rfloor}+mn^{4/5})$. Our method introduces only provably negligible error for Softmax attention. This work represents a significant step towards enabling efficient long-context processing in LLMs.
@@ 1 Introduction
In practical LLM applications, there are two scenarios for attention computation depending on the context length $n$ and query length $m$. The first case, $m=\Theta(1)$, represents the generation decoding based on the pre-computed Key Value Cache (KV), which stores the intermediate attention key and value matrices. The second case, $m=\Theta(n)$, represents the prompt prefilling before text generation or the cross-attention computation. However, in both cases, when the context window $n$ becomes larger, the running time will increase correspondingly, i.e., it will be linear and quadratic in $n$ for $m=\Theta(1)$ and $m=\Theta(n)$, respectively. Thus, reducing the running time of attention computations with long context input becomes essential to minimize response latency and increase throughput for LLMs.
In this work, we introduce novel methods to reduce the running time complexity for both cases, i.e., $m=\Theta(1)$ and $m=\Theta(n)$. We are inspired by the inherent sparsity within attention mechanisms. Numerous prior studies have highlighted the significant sparsity in the attention matrix [15, 4, 64, 86, 75]. This manifestation of sparsity in Softmax attention is that a large number of attention scores, i.e., $QK^{\top}$, concentrate on a small number of entries, which is known as “massive activation”. Due to this nature, Softmax attention can be accelerated by only calculating the entries that contain large attention scores, introducing negligible approximation errors [103, 49].
When talking about ReLU activation, one can easily accelerate the computation process by only calculating the entries activated by ReLU (since other non-activated entries will eventually be set to zero by ReLU). ReLU attention is another attention mechanism widely used, substituting the conventional Softmax activation function with ReLU. ReLU attention has demonstrated performance comparable to Softmax attention in various downstream tasks [91, 32] (see Section 2 for more details). We present the formal definition of ReLU attention as follows.
@@
To expedite the computation, the critical task is to identify the large/non-zero entries for Softmax/ReLU attention, respectively. In this work, We utilize the half-space reporting (HSR) data structure to tackle this problem. HSR was first introduced in [1] to address the half-space range reporting problem (More details can be found in Section 3.2).
In our framework, we define the half-space as the region where the attention scores (the inner products of key and query vectors) exceed some threshold. We leverage the HSR data structure’s ability to quickly answer range reporting to expedite the identification of non-zero entries within the ReLU attention matrix and large entries in Softmax attention. Consequently, we accelerate the computation for ReLU attention and expedite the computation of Softmax attention with negligible approximation error, resulting in a substantial reduction in computation time.
Then, we state our results under the generation decoding scenario ($m=\Theta(1)$) and prompt prefilling scenario ($m=\Theta(n)$). When $m=\Theta(1)$, we accelerate ReLU and Softmax attention computation time over the naive approach from $O(mn)$ to $O(mn^{4/5})$ with pre-processed KV cache (Algorithm 1). When $m=\Theta(n)$, we accelerate ReLU and Softmax attention computation time over the naive approach from $O(mn)$ to $O(mn^{1-1/\lfloor d/2\rfloor}+mn^{4/5})$ (Algorithm 2). Furthermore, Section 7 shows that the approximation error associated with Softmax attention utilizing “massive activated” entries only is small in practice, which is consistent with our theoretical analysis.
@@
  • To the best of our knowledge, this is the first work incorporating the HSR data structure with attention computation to reduce the running time complexity with the help of the sparsity within the attention mechanisms.
  • We provide rigorous theoretical proofs for reducing the computational time (1) for ReLU attention generation decoding from $O(mn)$ to $O(mn^{4/5})$ (Algorithm 1); (2) for ReLU attention prompt prefilling from $O(mn)$ to $O(mn^{1-1/\lfloor d/2\rfloor}+mn^{4/5})$ (Algorithm 2).
  • We achieve the same running time speed up for the conventional Softmax attention (Theorem 4.2, 5.2), and we give rigorous theoretical proofs to ensure that the approximation error remains negligible (Theorem 4.3). And we provide an empirical evaluation to prove our theoretical error analysis (Section 7).
@@
Comparison with Previous Works.
[25] uses an approximated nearest neighbors search, which requires that the data are well-conditioned, e.g., uniformly separated, while our algorithm supports an exact nearest neighbor search. Their nearest neighbors search may introduce large approximation errors, while the approximation error of our algorithm can be small. On the other hand, [5] uses low-rank approximation methods to accelerate the attention computation while they require bounded entries assumptions, which is not required in this work. The bounded entry assumption may not always be held in practical scenarios. Our work is based on practical observation that the attention matrix is sparse.
Roadmap.
Section 2 presents related work. Section 3 introduces essential concepts. Section 4 presents our main results, i.e., guarantees on run time reduction and approximation error. Section 5 introduces the extension of our method on prompt prefilling scenarios. Section 6 provides a summary of the techniques used in our proof. Section 7 provides our empirical evaluation for the approximation error on Softmax attention. Section 8 discusses the potential of extending our method. Section 9 concludes our algorithm and contributions.
@@ 2.1 Attention Acceleration for Long Context Input
A long context window is essential for transformer-based LLMs in many downstream tasks. However, due to the quadratic time complexity associated with attention mechanisms, transformers are usually hard to run inference efficiently. Numerous methods have been proposed to enhance the inference efficiency. One approach involves using alternative architectures as proxies for attention to support faster inference, such as Mamba [27, 22], PolySketchFormer [45], Hopfield Models [38, 90, 34, 97, 89, 30, 31] and Linearizing Transformers [99, 67]. Another line of research focuses on approximating attention matrix computation [5, 6, 7, 33, 101, 62, 70, 21, 60, 52, 29, 23, 58, 43, 16]. Nevertheless, these methods often rely on assumptions that may not be practical. For instance, some approaches use polynomial methods to approximate the exponential function, which requires all entries to be bounded by a small constant. However, our HSR-enhanced attention framework is designed based on practical observation and validated by empirical support. These advancements not only improve general model performance but also play a crucial role in enhancing in-context learning capabilities, where models leverage information from the immediate context to perform tasks without fine-tuning. We refer the readers to some other related works [57, 51, 53, 61, 62, 63, 80, 98, 77, 55, 83, 73, 81, 100, 84, 83, 12, 14, 18, 17, 78, 79, 19, 50, 44, 54, 92, 37, 36, 35, 94].
2.2 ReLU Attention
@@
ReLU attention employs the ReLU activation function in place of the traditional Softmax function for attention computation. Previous studies have highlighted the promising potential of ReLU attention in various domains. From the empirical side, [91] has demonstrated that incorporating ReLU as the activation function in vision transformers enhances performance on downstream tasks. [76] has shown that transformers equipped with ReLU attention outperform those with Softmax attention, particularly when dealing with large key-value memory in machine translation tasks. From the theoretical side, the scale-invariant property of ReLU attention [48] facilitates the scalability of transformer networks. Furthermore, [9, 26, 59] have shown that the inherent properties of ReLU attention contribute positively to the learning process of transformer models. Another key advantage is that the ReLU function effectively sets all negative values to zero, allowing us to bypass these non-contributory elements during attention computation and thereby reducing its running time. Omitting these zero and negative entries does not introduce any error into the final output of the ReLU attention mechanism.
2.3 Half-Space Reporting (HSR) Data Structure
@@
The HSR data structure, initially proposed by [1], was developed to address the half-space range reporting problem. The expedited range query capability inherent to HSR has been demonstrated to significantly enhance computational efficiency across a variety of tasks. Studies such as [41] and [11] have applied HSR to facilitate solving general linear programming (LP) problems. Another line of research has highlighted HSR’s potential in expediting the training process of contemporary neural networks [72, 28]. There is also a collection of research that concentrates on leveraging HSR for the advancement of solutions to geometric and graphical challenges [20, 40, 24].
@@ 3.1 Notations
We first introduce basic notations used in this paper. For any positive integer $n$, we use $[n]$ to denote set $\{1,2,\cdots,n\}$. We use $\operatorname*{\mathrm{Var}}[]$ to denote the variance. For two vectors $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $X_{i,j}$ to denote the $i$-row, $j$-th column of $X\in\mathbb{R}^{m\times n}$. We use $\|A\|_{\infty}$ to denote the $\ell_{\infty}$ norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{\infty}:=\max_{i\in[n],j\in[d]}|A_{i,j}|$.
@@ 3.2 Half-Space Reporting (HSR) Data Structure
Due to the limitation of space, we provide only the core corollary of the HSR data structure here. We refer the readers to Appendix B.4 for more details.
In [1], the author introduces a data structure named HSR to solve the half-space range reporting problem. The interface of the HSR data structure can be summarized as in Algorithm 3. Intuitively, the HSR data structure recursively partitions the set $S$ and organizes the points in a tree data structure. Then for a given query $(a,b)$, all $k$ points of $S$ with $\operatorname{sgn}(\langle a,x\rangle-b)\geq 0$ are reported quickly. Note that the query $(a,b)$ here defines the half-space $H$ in Definition B.10. We summarize the running time complexity of the HSR data structure as follows:
Corollary 3.1 (HSR data-structure time complexity [1], informal version of Corollary B.12).
@@
4 Main Results on Generation Decoding
In this section, we present our key findings regarding generation decoding, $m=\Theta(1)$, for both ReLU and Softmax attention mechanisms. We reduce the time complexity from a naive $O(mn)$ to $O(mn^{4/5})$. Our method only introduces a negligible approximation error for Softmax attention.
Figure 2: An outline of our principal algorithms. Top: Algorithm 1 for generation decoding is depicted, with the key matrix $K$ is fixed. During each inference step, the input query $Q$ interacts with the HSR data structure to get the activated indices set $\widetilde{S}_{i,j}$. Then, we can calculate the attention matrix according to $\widetilde{S}_{i,j}$. Bottom: Algorithm 2 for prompt prefilling is shown, where both the key matrix $K$ and the query matrix $Q$ are variable across iterations. Consequently, the HSR data structure must first be initialized with $K$, followed by querying it using $Q$. Finally, according to the activated entries set $\widetilde{S}_{i,j}$ reported by the HSR data structure, the attention matrix can be calculated. For more information, please refer to Remark 6.4.
Algorithm 1 Generation decoding
1:data structure GenerationDecoding $\triangleright$ Lemma 6.2